UT 20150119 LUCI2 AO Daytime Commissioning

Observer LUCI-Team: WSeifert
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator:
LBTO Support:

Plan for the day

Tasks #0 and #1 of SEspositos outline of AO tasks from LUCI2+FLAO daytime commissioning list

Summary

- Due to the fact that work could start already at 7.30am and continue until 3.15pm
more could be done than initially planned. Work was done on Tasks #0 (completed),
#1 (done for the moment), #4 (partly done) and #8 (partly done).

- Task #1: all modes (400) can only be closed at ROT=180...330deg for best NCPA during this run

- Remote readout hang up several times after a few frames. Modified version installed but still not perfect, needs more investigation

Details

All times are UT
Files are luci2.20150119.NNNN.fits

General checks

- Determination field center N375 to N30:
N375 #0001, N30 #0002
N375 center @ 1020/1020, upper right corner 1160/1170
N30 center 820/850, upper right corner 1960/1980
Goal for Retrorefl image on N375: 1090/1090

- Crossed filter: z & Pgam seems ok, around 5*10^-3 at 1.094\xB5m

- Remote readout: now 19s (1 image 3s exp time)

Task #0: Acq of reference source with retroreflector

- first on 3.75 camera approx centering
- on WFS improve IQ and do acq
-> diff lim image on 375 #0003
- centering & stage durection
1025/1035 #0003
1024/1043 stage y +1arcsec
1016/1043 stage x +1arcsec
1016/1060 stage y +2arcsec
1049/1059 stage x -4arcsec #0004
stage +x -> luci image -x
stage +y -> luci image +y
scale 1arcsec=8pix ok
- switch to N30
FeII #0005 2 spots, one close to focus
Brgam #0006 focus different due to chromatism of AO light source
- which is the good spot?
2nd spot comes from backside of EW reflection
bad spot re-imaged is closer to entrance window
-> move AO stage away from LUCI (minus)
-6mm #0007 bad (lower) spot in focus
+12mm both more out of focus -> back to orig
fine focus #0008 (looks like lab images!)
==> done

Task #1 (7.3.5 in Complan): Check for on-axsis vignetting on WFS vs. instrument rotator

- done by AO group 'internally'
-> all modes (400) can only be closed in ROT=180...330deg
-> during this run only in this range NCP can be well!
-> AI check the influence of ROT on NCP correction
==> done for the moment

Task #4 (7.3.7 in Complan): Best focus position for W unit Z stage

-> lower spot looks ok, in best focus @36mm
-> back to first (upper) spot
-> best focus @49mm
(difference in focus = 2*10mm -2*10/3mm=13mm)
--> upper spot is the 'good' one (as for IRTC and PISCES), the lower
one is a reflection from the backside of the EW
spot looks very similar to lab result; defocus by AO 1mm +-0.5mm
darks taken, 3s for AO: in DATA/20150120: luci2_20150119_0009/10/11
==> partially done

Task #8 (7.3.8 in Complan): Optical gain measurement

Check of applied aberation needed as preparation for ncp correction
==> partially done

Remote readout hang up several times after a few frames
-> Alexander investigated and installs modified version
now working better, but after 50 images again hang up
=> more work on needed

-- %USERSIG{JochenHeidt - 2015-01-20}%
Topic revision: r4 - 10 Mar 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback